# Materials Map Integrating Experimental and Computational Data through Graph-Based Machine Learning for Enhanced Materials Discovery

### v3

@@ Abstract
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of machine-learning and/or deep-learning techniques, researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods [6, 7, 8, 9, 10, 11, 12]. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project (MP) [13] and AFLOW [14] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [15]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications.
@@ 2 Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based representation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [15]. We use the dataset extracted by our previous study, described in Ref. [32] and briefly summarized in Appendix A. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [15], which provides a flexible environment for material property prediction. The model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models. MDL supports the construction of material maps using t-SNE with the data from the first dense layer, which is placed after GC layer as schematically described in Figure 1.
@@ 3.1 Material Map Generated Using MPNN Architecture
Let us first demonstrate a material map generated using the MPNN [28] architecture in Figure 2. The default hyperparameters of MDL [15] were used in the model training. The x and y axes represent the first and second t-SNE components, while the color of each point represents the predicted experimental value. In the figure, we find a clear trend where lower values concentrate in the lower region, while higher values appear in the upper region. The smooth color gradients reflect meaningful relationships between predicted values and structural features. Moreover, the fine structures visible in the map suggest that the model effectively captures structural features of materials. We also find two distinct branches, labeled BR1 and BR2 in Figure 2, extended laterally in the upper portion, featuring significant clusters of high values at their endpoints.
3.2 Contribution of GC Layer in MPNN Architecture for Map Construction
@@
In MDL, MPNN is defined by the Graph Convolutional (GC) layer configured by a neural network (NN) layer and a gated recurrent unit (GRU) layer. The NN layer enhances the model’s representational capacity, while the GRU layer improves learning efficiency through memory mechanisms. The GC layer is characterized by GC type, GC dimensions, and the repetition number of GC blocks (N). In this study, N was set to 4, which is the default hyperparameter of MDL [15], unless otherwise specified. To investigate the role of the GC layer, we compare maps generated by MDL with N = 1 and 10 in Figure 3. We find that the increase in N leads to tighter clustering of data points. To quantify the structures in the map, we evaluate the distributions of the nearest neighbor distances (NND) of each data point by Kernel Density Estimation (KDE) [33]. KDE is a non-parametric method for estimating smooth probability density functions from data points, and thus offers a more continuous and less biased distribution representation than histograms. Figure 3(c) summarizes the KDE spectra obtained with N = 1, 4 (default value for MDL), and 10. With increasing N, the peak intensity increases and the peak position shifts to the lower NND region, convincing the enhanced feature learning by the GC layer.
With increasing N, prediction accuracy and structural feature learning are enhanced, while the memory usage for computation increases dramatically especially when large datasets are used for the analysis [22]. This is demonstrated by the study using DeeperGATGNN, which allows modeling with large N [22]. Therefore, the use of MDL with MPNN architecture may be limited to cases where the dataset size is relatively small or when computing resources with sufficient memory size are available.
@@
In MDL, the MPNN architecture is characterized by NN and GRU blocks in the GC layer. To investigate the contributions of each block, we selectively controlled the training of each block, turning them ON or OFF, by modifying the MDL code. N was again set to 10 to enhance the contribution of each block. Figures 4(a)-4(d) illustrate the effects of disabling learning in either the NN or GRU blocks. When learning is disabled in either block, the points spread more diffusely; when both are disabled, the points distribute uniformly across the map. These trends are again clearly captured by the results of KDE, shown in Figure 4(e). By turning ON the training by either or both of the NN and GRU blocks, the peak intensity increases and the peak position shifts toward the lower NND region. These results convince us that the learning by the GC layer, composed of the NN and GRU blocks, plays a crucial role in the structural organization in the materials map.
3.3 Model Dependency of Material Maps
@@ 4 Discussion
4.1 Relationship between material property prediction and material map construction
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [32].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R2), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset are summarized in Table 1. While the MPNN produces the most structurally interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result leads us to conclude that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@ 4.1 Relationship between material property prediction and material map construction
4.2 Improvement of the interpretability of material maps
@@
4.3 Finding proper materials with material maps
@@
The material maps developed in this study have great potential to make materials development more efficient because materials with similar structures are likely to be synthesized and characterized using similar methods. For example, researchers working on high-entropy materials, which have significantly complex compositions and structures, require different knowledge and experimental equipment compared to those investigating ordinary materials. The interactive graph, shown in Figure 8, is a powerful tool for targeting specific material properties by referencing aspects such as composition, data names, and other material properties, enabling multifaceted and detailed data exploration. Using Python libraries such as Plotly, interactive graphs can be easily created, and when saved as HTML files, the data can be accessed at any time. The corresponding HTML file is included in summplementary data. In conclusion, the material maps, which organize materials data based on structural complexity, can assist researchers in screening candidate materials for experimental studies.
@@ 4.4 Prospects for future studies
Creation of materials maps requires substantial computational resources, especially for tasks such as dimensionality reduction by the t-SNE architecture and the deep-learning modeling defined in MDL. In this study, we employed the model and the hyperparameters given in MDL due to computational constraints. For wider and better applicability, fine-tuning of the model may be important. Recent advancements in high-performance computing (HPC) and cloud-based GPU/TPU clusters will help mitigate these computational challenges, enabling broader adoption of such methods.
In our study, we integrate computational and experimental data by predicting experimental values from compositions [32] by machine learning with experimental dataset of StarryData2 [31]. However, another approach—predicting stable structure from composition—is also conceivable. One should select the method based on the accuracies of the machine learning models, opting for the approach that achieves higher predictive accuracy. In this study, we employed our approach due to the high prediction accuracy of our model ( 0.9), predicting experimental values from compositions, and the technical difficulty in the prediction of the stable structure by composition [34].
Our study focused on the materials containing six selected elements (Sb, Te, Sn, Se, Bi, S), which demonstrated predictive capability for thermoelectric performance through the values [32]. Extending this model to encompass a wider range of elements and other physical parameters is left for future studies.
@@ 5 Conclusion
In this study, we developed and evaluated a novel approach for materials exploration by constructing a materials map that combines graph theory-based model and deep learning. A key feature of our work is the use of materials data from our previous research, which incorporated experimentally predicted values of material properties () along with structural information. The materials maps, generated using the MatDeepLearn (MDL) framework, were evaluated through distribution analysis of nearest-neighbor distances between data points by kernel density estimation. Results showed that among the graph theory-based models supported by MDL, the Message Passing Neural Network (MPNN) provided the most efficient structuring of material properties. Additionally, a systematic evaluation controlled training by the graph convolution layers of MDL demonstrated their significant contribution to the organization of the materials map.
@@ 9 Code availability
The code used for data analysis and simulated results is available from the corresponding authors upon reasonable request.
@@ A.1 Cleaning of Data in StarryData2
A.2 Machine Learning Modeling of Experimental Data
@@
This process resulted in a machine learning model with an R2 value of 0.85. Additional data cleaning, which exclude the data showing poor prediction accuracy, again improve the model, reading to Through the composition analysis of the selected materials, we selected Sb, Te, Sn, Se, Bi, and S as key elements for further investigation.
A.3 Machine Learning Modeling of Calculation Data
@@ Appendix B Correlation analysis of parameters
Appendix C Dimensionality Reduction for Materials Map Construction

### v4

@@ Abstract
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of data-science and machine-learning techniques, researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods [6, 7, 8, 9, 10, 11, 12]. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project (MP) [13] and AFLOW [14] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [15]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications.
@@ 2 Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based representation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [15]. We use the dataset extracted by our previous study, described in Ref. [32] and briefly summarized in Appendix A. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [15], which provides a flexible environment for material property prediction.
The machine learning model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models. The graph-based model is defined by the Graph Convolutional (GC) layer, which is characterized by GC type, GC dimensions, and the repetition number of GC blocks (N). In this study, N was set to 4, which is the default hyperparameter of MDL [15], unless otherwise specified. The GC layer for the MPNN architecture is configured by a neural network (NN) layer and a gated recurrent unit (GRU) layer. The NN layer enhances the model’s representational capacity, while the GRU layer improves learning efficiency through memory mechanisms.
MDL supports the construction of material maps using t-SNE. The data is extracted from the first dense layer, which is placed after GC layer as schematically described in Figure 1. This configuration allows the construction of material maps capturing both local and long-range structural correlations [15].
@@ 3.1 Material map generated using MPNN architecture
Let us first demonstrate in Figure 2 the material map generated using the MPNN architecture [28]. The default hyperparameters of MDL [15] were used in the model training. The x and y axes represent the first and second t-SNE components, while the color of each point represents the predicted experimental value. In the figure, we find a clear trend where lower values concentrate in the lower region, while higher values appear in the right- and left-upper region. The smooth color gradients reflect meaningful relationships between predicted values and structural features. Moreover, the fine structures visible in the map suggest that the model effectively captures structural features of materials. We also find two distinct branches, labeled BR1 and BR2 in Figure 2, extended laterally in the upper portion, featuring significant clusters of high values at their endpoints.
3.2 Contribution of GC layer for map construction
@@
To investigate the role of MPNN defined in the GC layer, we compare maps generated with N = 1 and 10 in Figure 3. We find that the increase in N leads to tighter clustering of data points. To quantify the structures of data points in the map, we evaluate the distribution of the nearest neighbor distances (NND) of each data point by Kernel Density Estimation (KDE) [33]. KDE is a non-parametric method for estimating smooth probability density functions from data points, and thus offers a more continuous and less biased distribution representation than histograms. Figure 3(c) summarizes the KDE spectra obtained with N = 1, 4 (default value for MDL), and 10. With increasing N, the peak intensity increases and the peak position shifts to the lower NND region, indicating the enhanced feature learning by the GC layer. It should be noted that the memory usage for computation increases dramatically with N especially when large datasets are used for the data analysis [22]. This is demonstrated by the study using DeeperGATGNN, which allows modeling with large N [22]. Therefore, the use of MDL with MPNN may be limited to cases where dataset size is small or when computing resources with sufficient memory size are available.
@@
In MDL, the MPNN architecture is characterized by NN and GRU blocks in the GC layer. To investigate the contributions of each block, we selectively controlled the training of each block, turning them ON or OFF, by modifying the MDL code. N was again set to 10 to enhance the contribution of each block. Figures 4(a)-(d) illustrate the effects of disabling learning in either the NN or GRU blocks. When learning is disabled in either block, the points spread more diffusely; when both are disabled, the points distribute uniformly across the map. These trends are again clearly captured by KDE as shown in Figure 4(e). By turning ON the training by either or both of the NN and GRU blocks, the peak intensity increases and the peak position shifts toward the lower NND region. These results convince us that the learning by the GC layer, composed of the NN and GRU blocks, plays a crucial role in the structural organization in the materials maps.
3.3 Model dependency of material map
@@ 4 Discussion
4.1 Relationship between material property prediction and material map
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [32].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R2), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset are summarized in Table 1. While MPNN produces the most structured and interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result means that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@ 4.1 Relationship between material property prediction and material map
4.2 Improvement of interpretability of material map
@@
4.3 Finding proper materials with material map
@@
The material maps developed in this study have great potential to make materials development more efficient because materials with similar structures are likely to be synthesized and characterized using similar methods. For example, researchers working on high-entropy materials [34], which have significantly complex compositions and structures, require different knowledge and experimental equipment compared to those investigating ordinary materials. The interactive graph, shown in Figure 8, is a powerful tool for targeting specific material properties by referencing aspects such as composition, data names, and other material properties, enabling multifaceted and detailed data exploration. Using Python libraries such as Plotly, interactive graphs can be easily created, and when saved as HTML files, the data can be accessed at any time. The corresponding HTML file is included in summplementary data. In conclusion, the material maps, which organize materials data based on structural complexity, can assist researchers in screening candidate materials for experimental studies.
@@ 4.4 Prospects for future studies
Creation of materials maps requires substantial computational resources, especially for tasks such as dimensionality reduction by the t-SNE architecture and the deep-learning modeling defined in MDL. In this study, we employed the model and the hyperparameters given in MDL due to computational constraints. For wider and better applicability, fine-tuning of the model may be important. Recent advancements in high-performance computing and GPU/TPU clusters will help mitigate these computational challenges, enabling broader adoption of such methods.
In our study, we integrate computational and experimental data by predicting experimental values from the compositions stored in the computational database [32]. However, another approach—predicting stable structure from composition—is also conceivable. One should select the method based on the accuracies of the machine learning models, opting for the approach that achieves higher predictive accuracy. In this study, we employed our approach due to the high prediction accuracy of our model ( 0.9), predicting experimental values from compositions, and the technical difficulty in the prediction of the stable structure by composition [35].
Our study focused on the thermoelectric properties of materials containing six selected elements (Sb, Te, Sn, Se, Bi, S), which demonstrated predictive capability by machine learning models trained by experimental data of values [32]. Extending this model to other physical properties and a wider range of elements is left for future studies.
@@ 5 Conclusion
In this study, we demonstrated the construction and evaluation of material maps generated by the MatDeepLearn (MDL) framework and the integrated datasets of computational and experimental studies. The map generated by MDL with MPNN architecture shows clear structures, representing the structural complexities of materials, with smooth color gradients reflecting the predicted experimental values. The structures in the maps are statistically evaluated using the distribution analysis of nearest neighbor distances (NND) for each data point through Kernel Density Estimation (KDE). Systematic analysis of material maps across different types, structures, and training configurations of graph-based models reveals the superior performance of MPNN for generating well-structured material maps that represent the structural similarities of materials.
Material maps generated using integrated datasets of experimental and computational studies have potential to bridge the gap between these approaches. Since materials with similar structures tend to be synthesized and evaluated with similar methods experimentally, material maps reflecting the similarity of material structures can support the targeting of promising materials by experimental researchers, and thus contribute to efficient and enhanced materials discovery.
@@ 9 Code availability
The codes used for data analysis and simulated results are available from the corresponding authors upon reasonable request.
@@ A.1 Cleaning of data in StarryData2
A.2 Machine learning modeling of experimental dataset
@@
This process resulted in a machine learning model with an R2 value of 0.85. Additional data cleaning, which exclude the data showing poor prediction accuracy, again improve the model, leading to Through the composition analysis of the selected materials, we selected Sb, Te, Sn, Se, Bi, and S as key elements for further investigation.
A.3 Machine learning modeling of calculation dataset
@@ Appendix B Correlation analysis of parameters
Appendix C Dimensionality reduction for material map construction